OpenStudy (anonymous):

Rewrite y = x^2 + 14x + 29 in general form

OpenStudy (anonymous):

Ignore the 29 temporarily and complete the square of the remaining part. Since \[\left(\frac{14}{2}\right)^2=49\] you have that \[y=x^2+14x+29=x^2+14x+49-49+29 = (x+7)^2-20\] Thus, the general form of your quadratic is \[y=(x+7)^2-20\]